# Cut-and-paste for impulsive gravitational waves with $\Lambda$: The mathematical analysis

### v1

@@ 6 Discussion
In this work we have studied the notorious discontinuous coordinate transformation (2.5) relating the distributional and the continuous metric commonly used to describe non-expanding impulsive gravitational waves propagating in (anti-)de Sitter space. Already in [39] it was shown that this transformation is geometrically given by the null generators of the (A)dS hyperboloid in a $5D$-description, which jump and are refracted due to the wave impulse. Here we have put this formal analysis on firm mathematical grounds using the nonlinear distributional analysis of the geodesics in these geometries provided in [45, 44]. More precisely, we have established that a careful geometric regularisation of the transformation leads to a generalized diffeomorphism in the sense of nonlinear distributional geometry. In this way we have also generalised the analysis of the far simpler $\Lambda=0$-case of [21]. We have schematically displayed our procedure in Figure 3.

### v2

@@ 6 Discussion
In this work we have studied the notorious discontinuous coordinate transformation (2.5) relating the distributional and the continuous metric commonly used to describe non-expanding impulsive gravitational waves propagating in (anti-)de Sitter space. Already in [39] it was shown that this transformation is geometrically given by the null generators of the (A)dS hyperboloid in a $5D$-description, which jump and are refracted due to the wave impulse. Here we have put this formal analysis on firm mathematical grounds using the nonlinear distributional analysis of the geodesics in these geometries provided in [45, 44]. More precisely, we have established that a careful geometric regularisation of the transformation leads to a generalised diffeomorphism in the sense of nonlinear distributional geometry. In this way we have also generalised the analysis of the far simpler $\Lambda=0$-case of [21]. We have schematically displayed our procedure in Figure 3.
@@ Acknowledgment
Data availability satement
Not applicable to this article because it is based on purely theoretical considerations, without using any datasets or other materials.
Conflict of interest statement
On behalf of all authors, the corresponding author (R.S.) states that there is no conflict of interest.